OpenStudy (anonymous):

MEDEL!!! 2x - y = 5 x + 3y = 7 What is the value of the system determinant?

HanAkoSolo (jamierox4ev3r):

you can solve this using elimination or substitution; does the problem specify which one you should use?

OpenStudy (anonymous):

multiply second equation with 2 and solve it

OpenStudy (anonymous):

x terms will be cancelled

HanAkoSolo (jamierox4ev3r):

that would be elimination @rathanreddy

HanAkoSolo (jamierox4ev3r):

2x=6y=14 is what is would become

OpenStudy (anonymous):

k sorry i ddn't see it

HanAkoSolo (jamierox4ev3r):

and actually @rathanreddy that is incorrect b/a it should be -2 to eliminate the other

OpenStudy (texaschic101):

he is looking for the determinant....this is not a square matrix....I do not know how to make your equations into a square matrix in order to find the determinant.

OpenStudy (anonymous):

ya when we solve we can change the signs

OpenStudy (anonymous):

 @Jamierox4ev3r

OpenStudy (texaschic101):

it has to be square, doesn't it....this is a 2x3 matrix....if it was a 2x2 or a 3x3 you could find the determinant...

OpenStudy (anonymous):

You can use substitution or elimination

OpenStudy (texaschic101):

wait...are you solving for the system of equations or are you looking for the determinant ?

OpenStudy (anonymous):

then use the way i told above you will get x and y

OpenStudy (anonymous):

in this one the determinant

OpenStudy (anonymous):

then you cant solve by elimination you should use determinant

OpenStudy (texaschic101):

I do not believe you can solve for the determinant if the matrix is not square...you have a 2x3 matrix...it is not square

OpenStudy (texaschic101):

@skullpatrol ....please take a look at this

OpenStudy (texaschic101):

@timo86m ...help

OpenStudy (texaschic101):

@johnweldon1993 ..help

OpenStudy (anonymous):

determinant = 2(3) - (-1)(1) = 6+1=7

OpenStudy (anonymous):

am i right?

OpenStudy (texaschic101):

I am not sure

OpenStudy (texaschic101):

@kropot72 ...please take a look at this

OpenStudy (texaschic101):

@satellite73 ....we need your assistance

OpenStudy (anonymous):

@satellite73

OpenStudy (johnweldon1993):

Hmm from what I can remember about linear equations and determinants....there are both an 'x' and a 'y' determinant....are you sure the question doesn't specify for either one?

OpenStudy (johnweldon1993):

Or there is just a coefficient determinant...I'm not sure which you need...if you need the coefficient determinant...than @DontLikeMathButOhWell would be correct here...

OpenStudy (texaschic101):

By any chance zachary, is this multiple choice ?....lol

OpenStudy (anonymous):

7 9 22

OpenStudy (texaschic101):

you might want to go with @DontLikeMathButOhWell ....go with 7

OpenStudy (johnweldon1993):

Yeah the 7 would look correct to me too lol

OpenStudy (texaschic101):

thanks john for coming and looking....good job Dontlikemath

OpenStudy (anonymous):

Yay :D, I feel accomplished

OpenStudy (johnweldon1993):

Lol good job @DontLikeMathButOhWell And no problem @texaschic101 :)